### v2

# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss first and second-order phase transitions and emerging new states including long-lived metastable states. We also discuss transitions between metastable states, chemical reactions and their products. These examples show that different values of FPCs would disable these transitions, and new phases as a result of these transitions would not emerge. This suggests that the current values of FPCs, by enabling transitions which give rise to new states, act to increase the phase space and entropy.
1 Introduction
Fundamental physical constants (FPCs), determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. Uzan (2003, 2011); Tiesinga et al. (2021); Ashcroft and Mermin (1976)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
Figure 1: Different scales of structure in the physical world including the range of typical sizes studied in condensed matter physics. Fairly recently, fundamental physical constants were shown to put lower and upper bounds on different condensed matter properties - properties that are already observed - including viscosity, diffusion and spin dynamics, electron and heat transport, thermal conductivity in insulators and conductors, the speed of sound, elastic moduli and velocity gradients that can be set up in cells using biochemical energy (see discussion in the Introduction). In section 3, we discuss a wider role of fundamental constants affecting observability and operation of entire phenomena and effects, including second-order structural and superconducting transitions, first-order melting transition, metastable transitions and chemical and other reactions.
Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). The “barcodes of ultimate reality” Barrow (2003), FPCs also set our Universe apart from those worlds where these constants might be different.
Understanding the values and origin of FPCs is one of the grandest challenges in modern science Barrow and Webb (2006). The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental Weinberg (1983). We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019)? This is done on the basis of our physical models. The bottom in Figure 1 involves length scales of particle physics. According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms. Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
The top in Figure 1 involves cosmological length scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life Barrow (2003); Adams (2019). The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. Adams (2019) discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
Compared to the above high-energy processes, the discussion of the role of FPCs at the scale of condensed matter has been scarce - this is the scale spanning 108-109 orders of magnitude in Figure 1 or about 20% of the logarithmic scale between the nuclei and the size of Universe. This has started to change fairly recently.
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a), velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b) and superconducting temperature Trachenko et al. . This has deepened our understanding of these properties and their physical models.
In this Perspective, we discuss a wider role of FPCs in condensed matter: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects. In the next section, we discuss why discussing FPCs at the condensed matter scale is important. This is followed by several examples in condensed matter physics showing how FPCs set observability and operation of phase transitions, the existence of different phases and states of matter emerging as a result of these transitions, including numerous metastable states. We also discuss chemical reactions and their products. These examples show that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This suggests that the current values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy. This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
2 Why filling the condensed matter gap is important
There are several reasons why understanding the role of FPCs at condensed matter length and energy scales and filling the condensed matter gap in Figure 1 is important. First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs Adams (2019). On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
Third, understanding the role of FPCs in condensed matter physics is an important enterprise on its own. It enriches the field itself and deepens understanding condensed matter phenomena at the most basic level. It also provides links and relationships to other areas of physics.
The fourth and final reason discussed in this Perspective is that if we better understand the operation of FPCs in condensed matter physics, we may find that the underlying principles are quite general and can be applied to other length and energy scales. This can lead to new insights related to the origin of FPCs and their values.
In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
3 Observability and operation of phase and other transitions. Existence of different phases and states
3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first is the set of FPCs. The second type are the environmental parameters such as, for example, temperature and pressure at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the temperature of the planet. In several examples in this section, we consider a certain range of environmental parameters (for example, the temperature range 102-103 K at typical planetary conditions in the Solar system) and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
We will be using and discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
$\beta=\frac{m_{p}}{m_{e}}$ (1)
and the fine-structure constant
$\alpha=\frac{e^{2}}{\hbar c}$ (2)
where $c$ is the speed of light in vacuum.
$\alpha$ and $\beta$ are of particular importance when discussing processes at different length scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter states, solids and liquids, are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
$a_{\rm B}=\frac{4\pi\epsilon_{0}\hbar^{2}}{m_{e}e^{2}}$ (3)
and
$E_{\rm R}=\frac{m_{e}e^{4}}{32\pi^{2}\epsilon_{0}^{2}\hbar^{2}}$ (4)
$a_{\rm B}$ is about 0.5 Å and gives a shorter range of typical interatomic separations condensed phases in the range 1-3 Å. $E_{\rm R}$ is about 13.6 eV and gives a larger value of cohesive energy in condensed phases in the typical range 1-10 eV.
3.2 Second-order phase transitions
Our first case study are the second-order phase structural transitions in solids. We consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. Let’s first consider $\beta$-cristobalite, the high-temperature phase emerging at the phase transition around 1800 K Hammonds et al. (1996); Dove (1997).
The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is Landau and Lifshitz (1970):
$S_{s}(T\gg\hbar\omega_{\rm D})=3N\left(\ln\frac{T}{\hbar\bar{\omega}}+1\right)$ (5)
where $\bar{\omega}$ is the average phonon frequency defined as $\bar{\omega}^{3N}=\prod\limits_{i}\omega_{i}$, $\omega_{i}$ are $3N$ phonon frequencies and $N$ is the number of atoms. Here and below, $k_{\rm B}=1$.
As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz. As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al. (1996); Dove (1997). This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq. (5).
The entropy (5) reflects the phase space available to $3N$ phonons in solids, and the increase of entropy of cristobalite is due to the phonon softening. However, the same entropy increase can be seen directly as the increase of the phase space of particle coordinates $q$ and momenta $p$, $\Delta\Gamma$. $\Delta\Gamma$ is:
$\Delta\Gamma=\frac{1}{(2\pi\hbar)^{3N}}\prod\limits_{i}\Delta p_{i}\Delta q_{i}$ (6)
Using $\langle\Delta p_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ and $\langle\Delta q_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ for each phonon mode in Eq. (6) and taking the relation between $\Delta\Gamma$ and $S$ as
$S=\ln\Delta\Gamma$ (7)
gives the entropy term $S\propto 3N\ln\frac{T}{\hbar\bar{\omega}}$ as in Eq. (5).
We see that smaller $\omega_{i}$ due to softer phonons in cristobalite increase $\langle\Delta p_{i}\rangle^{2}$ and $\langle\Delta q_{i}\rangle^{2}$. This increases $\Delta\Gamma$ and $S$ of cristobalite compared to quartz.
Cristobalite is a long-lived metastable phase of silica which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al. (1996); Dove (1997). According to Eq. (5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity). Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
As a more general case compared to pure metastable cristobalite, let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below the transition point where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence,
$S_{Q-C}>S_{Q}$ (8)
We therefore find that the entropy increase (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase. The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state at low $T$.
Without a phase transition operating in the considered range, the entropy would increase according to Eq. (5) within the single quartz phase but without the higher entropy contribution from the second metastable phase.
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ as
$\frac{\hbar\omega_{m}}{E}=\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (9)
where $m$ is the atom mass.
Setting $E=E_{\rm R}$ gives
$\omega_{m}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (10)
which has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
$\omega_{m}^{u}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\approx 3680\,\rm{K}$ (11)
$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1). As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (10)-(11). Note that this question is well-posed because it is possible to change $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
Larger $\omega_{m}$ implies larger $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) is no longer applicable because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is Landau and Lifshitz (1970)
$S_{s}(T\ll\hbar\omega_{\rm D})=\frac{4\pi^{4}}{5}N\left(\frac{T}{\hbar\omega_{\rm D}}\right)^{3}$ (12)
$S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$. This means that the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition. The phase transition becomes inoperative.
Therefore, the current values of fundamental constants, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy in Eq. (8).
A binary quartz-cristobalite mixture with increased entropy (8) is readily generalised to a collection of many systems existing in long-lived metastable states. Figure 2 illustrates this point. The increase of the phase space and entropy in these metastable states is generally expected because (a) high-temperature phases tend to be more symmetric Landau and Lifshitz (1970) with more open structure and have a larger number of soft phonons as a result Hammonds et al. (1996); Dove (1997) and (b) the increased phase space and entropy in the high-temperature phases become “frozen-in” in metastable states at low $T$.
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. The mixture of quartz and metastable cristobalite at low temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms). Phase transitions between different states enable sets of long-lived metastable systems with higher combined entropy at the same temperature. Shown minerals are for illustration only.
This case study gives us several insights. First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, can increase the phase space and entropy.
3.3 Superconducting transition
Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity. We have recently asked Trachenko et al. whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get? Fundamental limit here specifically means the limit in terms of FPCs.
Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (11) Trachenko et al. . $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$. This range of $\lambda$ corresponds to $C=0.1-0.2$. Using Eq. (11), this gives $T_{c}^{u}$ as Trachenko et al.
$T_{c}^{u}=C\omega_{m}^{u}=C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\approx 10^{2}-10^{3}\,\rm{K}$ (13)
Using Eq. (13), we ask how would the observability of superconductivity change were fundamental constants to take different values? Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved.
Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature. Consistent with Eq. (13) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides Flores-Livas et al. (2020); Duan et al. (2017); Pickard et al. (2019). This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above. We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
3.4 First-order phase transitions: solids, liquids, gases
Three basic states of matter, solids, liquids and gases, are represented by a phase diagram with three first-order phase transition lines as shown in Figure 3. In this section, we discuss how the observability of different states of matter and transitions between those states at our environmental conditions is related to FPCs.
Figure 3: Three states of matter, solids, liquids and gases, separated by three phase transition lines. The triple and critical points are shown.
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. Trachenko (2024) for the functional form of pressure-temperature melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds. This cohesive energy is set by FPCs in Eq. (4). Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq. (9)). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs. This gives the following role of FPCs in setting the states of matter.
FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions. Melting and boiling phase transitions would be unobservable. Conversely, FPCs resulting in low cohesive energies and melting temperatures imply that matter would exist in the gas state only.
We see that, similarly to second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life. First, H2O can exist as liquid water at our typical planetary conditions. Second, carbon-based substances can exist as bonded solidlike structures. Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures would not exist as cohesive states but as the gas states only.
We note that classical entropies of solids and gases contain FPCs explicitly. Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ Landau and Lifshitz (1970), $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq. (10) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
$S_{s}\propto\ln\frac{\hbar^{2}m_{p}^{\frac{1}{2}}}{e^{4}m_{e}^{\frac{3}{2}}}$ (14)
The entropy of the classical gas is Landau and Lifshitz (1970)
$S_{g}=\frac{3}{2}N\left(\ln T+1\right)+N\left(\ln v+\ln\left(\frac{m}{2\pi\hbar^{2}}\right)^{\frac{3}{2}}+1\right)$ (15)
where $v$ is the volume per atom and $m$ is the atom mass.
With $v\propto a_{\rm B}^{3}$ and $m\propto m_{p}$,
$S_{g}\propto\ln\frac{\hbar^{3}m_{p}^{\frac{3}{2}}}{e^{6}m_{e}^{3}}$ (16)
The same FPCs interestingly feature in both numerators and denominators in Eqs. (14) and (16). $S_{g}$ in Eq. (16) is more sensitive to FPCs as compared to $S_{s}$ in Eq. (14): larger $\hbar$ and smaller $m_{e}$ and $e$ increase $S_{g}$ faster. The same variation - larger $\hbar$ and smaller $m_{e}$ and $e$ also decrease the cohesive energy in Eq. (4), favouring gases over solids. Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective.
Differently to solid and gas entropies we considered, no closed equation for the liquid entropy exists which corresponds to experimental properties of real liquids such as $c_{v}=3$ close to melting ($c_{v}$ is the constant-volume specific heat). This is due to the fundamental no-small parameter problem of the liquid theory (see Ref. Trachenko (2024) for a quick review and Refs. Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion). Existing models such hard-sphere and Van der Waals models and their extensions give $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically. However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: energy of the system is the sum of energies of its excitations Lifshitz and Pitaevskii (2006). In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024). The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.
3.5 Metastable states and chemical reactions
In previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. Brazhkin (2006) for review). When discussing phase transitions and phase diagrams in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006). Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al. .
Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs. As in the equilibrium case, these transitions correspond to the free energy crossover. Hence, similarly to what we saw for equilibrium transitions in earlier sections, the observability and operation of transitions between different metastable states and hence the existence of numerous metastable phases is set by FPCs.
Similarly to what we saw in Section 3.2, the entropy increases due to new metastable states emerging as a result of these transitions because the states formed at higher temperature tend to be more symmetric with more open structure and have larger phase space as a result (this larger phase space and higher entropy become frozen in the low-temperature metastable phase).
The same assertion of observability of transitions and different phases applies to chemical reactions and reactions products. Chemical reactions involve changes of chemical bonding whose characteristic energies are set by FPCs. Hence these reactions are governed by the interplay of characteristic cohesive energies set by FPCs and environmental parameters. For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be unobservable at given environmental parameters.
4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019). A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007). FPCs would then be calculable from this theory.
In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the entropy. Long-lived metastable states were important in this discussion because at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy. However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)). These long-lived metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars. In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and long-lived metastable phases emerging, with the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of fundamental constants act to increase the entropy. This insight is general and can be applied to different levels of structure in the Universe in Figure 1. This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism (see Ref. Trachenko (2023d) for the review of related earlier discussions) where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability (fluctuations) of FPCs, analogous to genetic mutation in biological evolution, is currently unknown, although several mechanisms were discussed (see, e.g., Refs. Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013)).
Cyclic or oscillatory models of the Universe (see, e.g., Refs. Adams and Laughlin (1997); Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013); Trachenko (2017) and Ref. Kragh (2009) for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe in Ref. Trachenko (2017) at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space or close to it. Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis in stars and so on). The evolutionary mechanism (with the entropy increase as a selection principle) would imply that structures observed in each cycle are related to the magnitude of fluctuations of FPCs. Large fluctuations would destroy the intricate network of constraints on the values of FPCs needed for different structures in the Universe to exist Adams (2019), resulting in the scarcity of structures and smaller phase space. Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles. There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (this selection applies to FPCs and FPC-enabled structures existing at the time). This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence. This process involves maximising a property (entropy) as a function of FPCs but within many constraints on their values required to keep different structures and their phase space intact. This invites an input from the optimisation theory.
One example illustrating some of the discussion in the previous paragraph, but which needs further study, is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
$S_{BH}=4\pi\frac{G}{\hbar c}M_{BH}^{2}$ (17)
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on. Given the enormity of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
5 Discussion and summary
Approaching the challenge to explain fundamental constants started with understanding the role these constants play. This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties. This has enriched condensed matter physics and deepened our understanding of its key phenomena.
Here, we discussed a wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena. We gave examples of operation of first- and second-order phase transitions and emerging new phases, existence of numerous metastable phases, chemical reactions, their products and so on. We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This suggests that the values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
This list of effects and phenomena in condensed matter can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. Adams (2019) for a review of these challenges).
If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe. The insight that FPCs act to increase the entropy is one such example. This and similar insights can provide useful links between different areas of physics and open up new discussions.
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and superconductivity Trachenko et al. . Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.
I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.