### v3

# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. We also discuss metastable states, transitions between them, chemical reactions and their products. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs. We show that the new states emerging as a result of various transitions increase the phase space and entropy. Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge. Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy.
1 Introduction
Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. Uzan (2003, 2011); Tiesinga et al. (2021); Ashcroft and Mermin (1976)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
Figure 1: Different scales of structure in the physical world including the range of typical sizes studied in condensed matter physics. Fairly recently, fundamental physical constants were shown to put lower and upper bounds on different condensed matter properties - properties that are already observed - including viscosity, diffusion and spin dynamics, electron and heat transport, thermal conductivity in insulators and conductors, the speed of sound, elastic moduli and velocity gradients that can be set up in cells using biochemical energy (see discussion in the Introduction). In section 3, we discuss a wider role of fundamental constants affecting observability and operation of entire phenomena and effects, including second-order structural and superconducting transitions, first-order melting transition, metastable transitions and chemical and other reactions.
Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). The “barcodes of ultimate reality” Barrow (2003), FPCs also set our Universe apart from those worlds where these constants might be different.
Understanding the values and origin of FPCs is one of the grandest challenges in modern science Barrow and Webb (2006). The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental Weinberg (1983). We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019)? This is done on the basis of our physical models. The bottom in Figure 1 involves length scales of particle physics. According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms. Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life Barrow (2003); Adams (2019). The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. Adams (2019) discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
Compared to the above high-energy processes, the discussion of the role of FPCs at the scale of condensed matter has been scarce - this is the scale spanning 108-109 orders of magnitude in Figure 1 or about 20% of the logarithmic scale between the nuclei and the size of Universe. This has started to change fairly recently.
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a), velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b) and superconducting temperature Trachenko et al. . This has deepened our understanding of these properties and their physical models.
In this Perspective, we discuss a wider role of FPCs: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects. In the next section, we discuss why discussing FPCs at the condensed matter scale is important. This is followed by considering second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. This discussion shows how FPCs set the observability and operation of phase transitions and hence the existence of different states emerging as a result of these transitions, including metastable states. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature in, for example, structural phase transitions, can be calculated from FPCs.
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states due to these transitions would not emerge. Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy. This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
2 Why filling the condensed matter gap is important
There are several reasons why understanding the role of FPCs at condensed matter length and energy scales and filling the condensed matter gap in Figure 1 is important. First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs Adams (2019). On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
Third, understanding the role of FPCs in condensed matter physics is an important enterprise on its own. It enriches the field itself and deepens understanding condensed matter phenomena at the most basic level. It also provides links and relationships to other areas of physics.
The fourth and final reason discussed in this Perspective is that if we better understand the operation of FPCs in condensed matter physics, we may find that the underlying principles are quite general and can be applied to other length and energy scales. This can lead to new insights related to the origin of FPCs and their values.
In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
3 Observability and operation of phase and other transitions. Emergence of new states and entropy
3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first type are FPCs. The second type are the environmental parameters such as temperature and pressure at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the planet’s temperature. In several examples in this section, we consider a certain range of environmental parameters ($\Delta T$,$\Delta P$) around, for example, temperature (102-103 K) and pressure range at typical planetary conditions in the Solar system and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
We will be using and discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
$\beta=\frac{m_{p}}{m_{e}}$ (1)
and the fine-structure constant
$\alpha=\frac{e^{2}}{\hbar c}$ (2)
where $c$ is the speed of light in vacuum.
$\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
$a_{\rm B}=\frac{4\pi\epsilon_{0}\hbar^{2}}{m_{e}e^{2}}$ (3)
and
$E_{\rm R}=\frac{m_{e}e^{4}}{32\pi^{2}\epsilon_{0}^{2}\hbar^{2}}$ (4)
$a_{\rm B}\approx 0.5$ and is on the short side of interatomic separations in condensed phases in the typical range 1-3 Å. $E_{\rm R}\approx$ 13.6 eV and is on the large side of cohesive energies in condensed phases in the range 1-10 eV.
3.2 Second-order phase transitions
3.2.1 Enropy increase due to a phase transition
Our first case study are second-order phase structural transitions in solids. Let’s consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition Hammonds et al. (1996); Dove (1997).
The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is Landau and Lifshitz (1970):
$S_{s}(T\gg\hbar\omega_{\rm D})=3N\left(\ln\frac{T}{\hbar\bar{\omega}}+1\right)$ (5)
where $\bar{\omega}$ is the average phonon frequency defined as $\bar{\omega}^{3N}=\prod\limits_{i}\omega_{i}$, $\omega_{i}$ are $3N$ phonon frequencies and $N$ is the number of atoms. Here and below, $k_{\rm B}=1$.
As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz. As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al. (1996); Dove (1997). This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq. (5).
The entropy (5) reflects the phase space available to $3N$ phonons in solids, and the increase of entropy of cristobalite is due to the phonon softening. However, the same entropy increase can be seen directly as the increase of the phase space of particle coordinates $q$ and momenta $p$, $\Delta\Gamma$. $\Delta\Gamma$ is:
$\Delta\Gamma=\frac{1}{(2\pi\hbar)^{3N}}\prod\limits_{i}\Delta p_{i}\Delta q_{i}$ (6)
Using $\langle\Delta p_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ and $\langle\Delta q_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ for each phonon mode in Eq. (6) and
$S=\ln\Delta\Gamma$ (7)
gives the entropy term $S\propto 3N\ln\frac{T}{\hbar\bar{\omega}}$ as in Eq. (5).
We see that smaller $\omega_{i}$ due to softer phonons in cristobalite increase $\langle\Delta p_{i}\rangle^{2}$ and $\langle\Delta q_{i}\rangle^{2}$. This increases $\Delta\Gamma$ and $S$ of cristobalite compared to quartz.
Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$. Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are Dove (1997, 2003):
$\tilde{\omega}_{i}^{2}=\omega_{i}^{2}+\frac{1}{2}\delta_{i}\eta^{2}$ (8)
where $\omega_{i}$ are harmonic frequencies and $\delta_{i}$ are the anharmonicity coefficients.
In the high-temperature phase above the phase transition temperature $T_{c}$, $\eta=0$. Phonon frequencies become smaller as a result, increasing the entropy.
Importantly, the entropy increase is ultimately related to a phase transition which gives rise to the second high-entropy phase. Without a phase transition operating, the entropy would increase according to Eq. (5) within the single low-temperature phase only but without larger entropy coming from the second high-temperature phase.
The entropy increase due to the phase transition can take place both above and below $T_{c}$. For reversible phase transition where the low-temperature phase is always recovered below $T_{c}$, the entropy increase takes place above $T_{c}$. In our model quartz-cristobalite example, this gives
$S_{C}(T>T_{c})>S_{Q}(T>T_{c})$ (9)
where $S_{C}(T>T_{c})$ is the entropy of the high-temperature cristobalite phase above $T_{c}$ and $S_{Q}(T>T_{c})$ is the entropy that quartz would have above $T_{c}$ without the phase transition operating. More generally, (9) applies to any two low and high-temperature phases.
The entropy increase can also take place below $T_{c}$ if the high-temperature phase becomes a long-lived metastable state below $T_{c}$. In our model quartz-cristobalite example, the phase transition is a reconstructive phase transition with large energy barriers separating the two phases. As a result, cristobalite is a long-lived metastable phase which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al. (1996); Dove (1997). According to Eq. (5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity). Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence,
$S_{Q-C}(T<T_{c})>S_{Q}(T<T_{c})$ (10)
The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state below $T_{c}$.
The quartz-cristobalite example is readily generalised to many systems where the combined entropy increases due to phase transitions. Figure 2 illustrates this point.
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. In a set of solids, phase transitions result in larger phase space and higher entropy. Shown minerals are for illustration only.
3.2.2 Observability of phase transitions. Entropy without a phase transition
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ Trachenko and Brazhkin (2020)
$\frac{\hbar\omega_{m}}{E}=\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (11)
where $m$ is the atom mass. Setting $E=E_{\rm R}$ gives
$\omega_{m}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (12)
$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
$\omega_{m}^{u}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\approx 3680\,\rm{K}$ (13)
$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1). As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (12)-(13). Note that this question is well-posed because it is possible to vary $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is Landau and Lifshitz (1970)
$S_{s}(T\ll\hbar\omega_{\rm D})=\frac{4\pi^{4}}{5}N\left(\frac{T}{\hbar\omega_{\rm D}}\right)^{3}$ (14)
$S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$. Then the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition. The phase transition becomes inoperative.
The observability of phase transitions can also be discussed on the basis of $T_{c}$. In phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), the order of magnitude of $T_{c}$ is set by the energy of interaction between particles Dove (1997). This energy is on the order of the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$. Using Eq. (12), this gives
$T_{c}\approx\hbar{\omega_{m}}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{2}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (15)
To estimate $T_{c}$, we write $m=Am_{p}$, where $A$ is the atomic mass. Using Eqs. (12) and (13), $T_{c}$ in Eq. (15) is
$T_{c}=\frac{\hbar\omega_{m}^{u}}{A^{\frac{1}{2}}}$ (16)
$A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1 to 15, with an average value of 8. Using this value and $\omega_{m}^{u}=3680$ K in Eq. (13), Eq. (16) gives $T_{c}\approx$ 500 K. This is close to $T_{c}$ calculated from typical parameters of the anharmonic phonon model Dove (1993) and is of the right order of magnitude of $T_{c}$ in the range $10^{2}-10^{3}$ K typically observed.
We see that the order of magnitude of $T_{c}$ (15) is set by FPCs. Were different values of FPCs to give $T_{c}$ on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions. This is consistent with our earlier argument involving small entropy in Eq. (14) due to large $\omega_{\rm D}$.
From this and previous sections, we conclude that the current values of FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy.
3.2.3 Summary
The discussion in Section 3.2 gives us several insights. First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, increase the phase space and entropy.
3.3 Superconducting transition
Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity. We have recently asked Trachenko et al. whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get? Fundamental limit here specifically means the limit in terms of FPCs.
Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (13) Trachenko et al. . $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$. This range of $\lambda$ corresponds to $C=0.1-0.2$. Using Eq. (13), this gives $T_{c}^{u}$ as Trachenko et al.
$T_{c}^{u}=C\omega_{m}^{u}=C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\approx 10^{2}-10^{3}\,\rm{K}$ (17)
Using Eq. (17), we ask how would the observability of superconductivity change were fundamental constants to take different values? Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved.
Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature. Consistent with Eq. (17) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides Flores-Livas et al. (2020); Duan et al. (2017); Pickard et al. (2019). This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above. We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
3.4 First-order phase transitions: solids, liquids, gases
Three basic states of matter, solids, liquids and gases, are represented by a phase diagram with three first-order phase transition lines as shown in Figure 3. In this section, we discuss how the observability of different states of matter and transitions between those states at our environmental conditions is related to FPCs.
Figure 3: Three states of matter, solids, liquids and gases, separated by three phase transition lines. The triple and critical points are shown.
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. Trachenko (2024) for the functional form of pressure-temperature melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds. This cohesive energy is set by FPCs in Eq. (4). Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq. (11)). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs. This gives the following role of FPCs in setting the states of matter.
FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions. Melting and boiling phase transitions would be unobservable. Conversely, FPCs resulting in small cohesive energies and hence low melting and boiling temperatures imply that matter would exist in the gas state only.
We see that, similarly to second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life. First, H2O can exist as liquid water at our typical planetary conditions. Second, carbon-based substances can exist as bonded solidlike structures. Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures would not exist as cohesive states but as the gas states only.
We note that classical entropies of solids and gases contain FPCs explicitly. Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ Landau and Lifshitz (1970), $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq. (12) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
$S_{s}\propto\ln\frac{\hbar^{2}m_{p}^{\frac{1}{2}}}{e^{4}m_{e}^{\frac{3}{2}}}$ (18)
The entropy of the classical gas is Landau and Lifshitz (1970)
$S_{g}=\frac{3}{2}N\left(\ln T+1\right)+N\left(\ln v+\ln\left(\frac{m}{2\pi\hbar^{2}}\right)^{\frac{3}{2}}+1\right)$ (19)
where $v$ is the volume per atom and $m$ is the atom mass.
With $v\propto a_{\rm B}^{3}$ and $m\propto m_{p}$,
$S_{g}\propto\ln\frac{\hbar^{3}m_{p}^{\frac{3}{2}}}{e^{6}m_{e}^{3}}$ (20)
The same FPCs interestingly feature in both numerators and denominators in Eqs. (18) and (20). $S_{g}$ in Eq. (20) is more sensitive to FPCs as compared to $S_{s}$ in Eq. (18): larger $\hbar$ and smaller $m_{e}$ and $e$ increase $S_{g}$ faster. The same variation - larger $\hbar$ and smaller $m_{e}$ and $e$ also decrease the cohesive energy in Eq. (4), favouring gases over solids. Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective.
Similarly to what we saw in Section 3.2 for the second-order transitions, the first-order phase transitions and the emerging high-temperature gas phase increase the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat). Similarly to the second-order transitions, the first-order transitions are enabled by FPCs as discussed earlier in this section.
Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref. Trachenko (2024) for a quick review and Refs. Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting Trachenko and Brazhkin (2016); Trachenko (2023c). However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations Lifshitz and Pitaevskii (2006). In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024). The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.
3.5 Metastable states and chemical reactions
In the previous three sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. Brazhkin (2006) for review). When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006). Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al. .
Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs. As in the equilibrium case, these transitions correspond to the free energy crossover. Hence, similarly to what we saw for equilibrium transitions in earlier sections, the observability and operation of transitions between different metastable states and hence the existence of numerous metastable phases is set by FPCs.
Similarly to what we saw in Section 3.2, the entropy increases due to new metastable states emerging as a result of these transitions because the states formed at higher temperature tend to be more symmetric with more open structure and have larger phase space as a result (this larger phase space and higher entropy become frozen in the low-temperature metastable phase).
The same assertion of observability of transitions and different phases applies to chemical reactions and reactions products. Chemical reactions involve changes of chemical bonding whose characteristic energies are set by FPCs. Hence these reactions are governed by the interplay of characteristic cohesive energies set by FPCs and environmental parameters. For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be unobservable at given environmental parameters.
4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019). A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007). FPCs would then be calculable from this theory.
In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy. We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of fundamental constants act to increase the entropy. This insight is general and can be applied to different levels of structure in the Universe in Figure 1. This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
We recall that the discussion in Section 3 included long-lived metastable states. We saw that at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy. As recently reviewed Brazhkin (2024), metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)). These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars. In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, and its operation are unknown, although several proposals were discussed (see, e.g. Refs. Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013) and Ref. Trachenko (2023d) for review).
Cyclic or oscillatory models of the Universe (see, e.g., Refs. Adams and Laughlin (1997); Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013); Trachenko (2017) and Ref. Kragh (2009) for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe in Ref. Trachenko (2017) at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space and entropy or close to it. Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis in stars and so on). The evolutionary mechanism (with the entropy increase as a selection principle) would imply that structures observed in each cycle are related to the magnitude of fluctuations of FPCs. Large fluctuations would destroy the intricate network of constraints on the values of FPCs needed for different structures in the Universe to exist Adams (2019), resulting in the scarcity of structures and smaller phase space. Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles. There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (similarly to biological evolution, this selection applies to FPCs and FPC-enabled structures existing at the time). This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence. This process involves maximising a property (entropy) as a function of FPCs but within many constraints on their values required to keep different structures and their phase space intact. This invites an input from the optimisation theory.
An example related to this discussion and involving constraints is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
$S_{BH}=4\pi\frac{G}{\hbar c}M_{BH}^{2}$ (21)
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), breaking many constraints required for key processes of structure formation and disrupting these processes Barrow (2003); Adams (2019). This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs). We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on. Given the scale of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
5 Discussion and summary
Approaching the challenge to explain fundamental constants started with understanding the role these constants play. This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties. This has enriched condensed matter physics and deepened our understanding of its key phenomena.
Here, we discussed a wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena. We gave examples of operation of first- and second-order phase transitions and emerging new phases, existence of numerous metastable phases, chemical reactions, their products and so on. We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This suggests that the values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
This list of effects and phenomena in condensed matter can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. Adams (2019) for a review of these challenges).
If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe. The insight that FPCs act to increase the entropy is one such example. This and similar insights can provide useful links between different areas of physics and open up new discussions.
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and superconductivity Trachenko et al. . Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.
I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.